# Contrastive Learning for Robust Representations of Neutrino Data

### v1

@@ I Introduction
The rest of the paper is organised as follows: Section II provides an overview of contrastive learning and the state-of-the-art, while Section III describes the methodology followed. Section IV describes the datasets used, Section VI presents the experimental results, and Section VII summarises the main findings and outlines directions for future work.
@@ II Contrastive learning
More recently, contrastive learning has regained significant attention as a powerful approach for domain adaptation by leveraging data augmentations and similarities between data points to learn robust, transferable representations [27, 28, 29]. By encouraging the alignment of similar samples while separating dissimilar ones, contrastive learning is particularly well-suited for addressing discrepancies between domains, such as the data-MC mismatch in high-energy and neutrino physics.
@@ IV.2 Segmented Scintillator Cube Detector
The dataset consists of simulated data for a three-dimensional fine-grain plastic scintillator detector with a size of $2\times 2\times 2~{}\mathrm{m^{3}}$. The detector is divided into $200\times 200\times 200$ cubic voxels of $1~{}\mathrm{cm^{3}}$ each, where each voxel acts as a sensor for scintillation light produced by traversing particles. Single-charged particles (protons, pions, muons, and electrons) with isotropic directions and uniform momentum distributions are simulated within the detector volume. Detector response simulation accounts for crosstalk, the leakage of scintillation light from one voxel into adjacent voxels, modelled with a 3% leakage probability per face. This publicly available dataset can be accessed at: https://doi.org/10.5281/zenodo.10998285.
@@ V Augmentations
We use a series of augmentations during the training of the contrastive learning models, as well as when training a baseline classifier with augmentations. We apply the same augmentations to both datasets. These datasets were augmented using a series of transformations designed to introduce spatial and energy variations, simulating realistic conditions and improving model robustness. During training, for both the contrastive models and the classifier with augmentations, we randomly sample with replacement a series of three augmentations from a full set of augmentation transformations. The augmentations applied include: a rotation transformation that randomly rotates the spatial coordinates; a masking operation that randomly drops a fraction of the input points, effectively simulating partial data loss; an energy scaling augmentation that introduces voxel-wise energy shifts by scaling the feature values with random noise sampled from a normal distribution; a spatial translation transformation that displaces the input points by a random offset proportional to a predefined scale factor; and the identity transformation. These augmentations ensured compatibility with batched processing and scalable data handling.
We also experimented with augmentation via re-simulation of the detector response for the LArTPC dataset. We re-simulated events at the larnd-sim level, altering parameters related to the electric field, transverse and longitudinal diffusion and electron lifetime. However, when using these augmentations with realistic simulation parameter ranges, we found the augmentations to be too weak. In line with previous work (making the case for using strong augmentations) [9], our trained contrastive model using purely re-simulation was unable to produce any useful features for downstream tasks. We also tried using re-simulation augmentations to create additional augmentations to complement the ones described in the previous paragraph. This resulted in marginal gains even when evaluated on samples that had undergone re-simulation with the same altered parameters as used in the augmentations. Given the significant computational requirement, we advocate against the use of re-simulation for augmentations in this use case.
@@ VI Experimental results
Figure 3: Bar chart of test accuracy of our methods - unsupervised and supervised contrastive learning, compared to DANN and a classifier with training data augmentations. Nominal indicates the test accuracy of the models using the same detector simulation parameters used in training (i.e., no distribution shift). During testing, these methods are evaluated on various data throws - shifts in detector simulation parameters chosen randomly as shown in Table 1.
Figure 4: Top: Accuracy of our proposed method - supervised contrastive learning compared to classifier with and without augmentations on the scintillator detector dataset. Bottom: Relative change from nominal (training) level of cross-talk to the current (testing) cross-talk level. Nominal cross-talk is the level that was used to train the model. The models are evaluated on increments of 0.05 cross-talk ranging from no cross-talk to the maximum level of 1.00.
@@
We evaluated the robustness of four different approaches: a classifier with augmentations, unsupervised contrastive learning, supervised contrastive learning, and DANN. Under nominal conditions, the classifier with augmentations achieved the highest performance at 87.3%, followed closely by our proposed method based on supervised contrastive learning (87.0%). The unsupervised contrastive learning approach resulted in a 3% drop in accuracy due to its inability to utilise the label information present in the dataset during the pre-training phase. The DANN approach led to a substantial performance decline, reaching only 74%, primarily due to the challenges in the adversarial training and the limited availability of diverse training examples for each throw.
@@
In Figure 4, we summarise the results of our proposed method on the scintillator detector dataset. Our method, based on supervised contrastive learning, is compared against a classifier using the same architecture. The classifier is evaluated under two conditions: (i) without augmentations during training, and (ii) with the same augmentations applied in the contrastive learning pre-training setup. We generate datasets with different levels of cross-talk by masking out fractions of the detector response induced by cross-talk in the simulation. We train separate models for each cross-talk level and evaluate these models across all cross-talk levels. The supervised contrastive learning model consistently outperforms the baseline classifier models, both with and without augmentations. Moreover, the contrastive setup maintains superior accuracy as the cross-talk level varies, highlighting that the extracted features are more dependent on the fundamental physical properties of the particle interactions generating the events. In contrast, the classifier-based approach appears less robust, potentially indicating that its features are more sensitive to specific properties of the detector simulation.
These results strongly suggest that contrastive learning approaches, particularly the supervised variant, provide enhanced stability against detector simulation variations compared to traditional deep learning classification methods. The study of both unsupervised and supervised approaches indicates that it is the contrastive learning approach, rather than the ability to pre-train on an unlabelled target distribution as in the unsupervised case, that results in representations robust to distribution shifts. This increased stability is particularly valuable in high-energy physics and related domains, where minor variations in detector conditions can significantly impact model performance. By leveraging contrastive learning, we can ensure that learned representations generalise more effectively across varying conditions, thereby reducing the risk of model performance degradation due to distribution shifts.

### v2

@@ I Introduction
The rest of the paper is organised as follows: Section II provides an overview of contrastive learning and the state-of-the-art, while Section III describes the methodology followed. Section IV describes the datasets used, and Section V details the augmentations applied. Section VI presents the experimental results, and Section VII summarises the main findings and outlines directions for future work.
@@ II Contrastive learning
More recently, contrastive learning has regained significant attention as a powerful approach for domain adaptation by leveraging data augmentations and similarities between data points to learn robust, transferable representations [27, 28, 29]. By encouraging the alignment of similar samples while separating dissimilar ones, contrastive learning is particularly well-suited for addressing discrepancies between domains, such as the data-MC mismatch in high-energy and neutrino physics. Recent methods like RS3L [30] use re-simulation to generate physics-driven augmentations, improving model robustness. Large-scale pretraining with contrastive learning has also been explored for collider data, demonstrating gains in adaptability and efficiency [31].
@@ IV.2 Segmented Scintillator Cube Detector
The dataset consists of simulated data for a three-dimensional fine-grain plastic scintillator detector with a size of $2\times 2\times 2~{}\mathrm{m^{3}}$. The detector is divided into $200\times 200\times 200$ cubic voxels of $1~{}\mathrm{cm^{3}}$ each, where each voxel acts as a sensor for scintillation light produced by traversing particles. Single-charged particles (protons, pions, muons, and electrons) with isotropic directions and uniform momentum distributions are simulated within the detector volume. Detector response simulation accounts for crosstalk, the leakage of scintillation light from one voxel into adjacent voxels, modelled with a 3% leakage probability per face. Each event in the dataset $x$ is a tensor $v\in\mathbb{R}^{N}$ containing the number of detected photoelectrons per voxel (cube) and their corresponding coordinates $c\in\mathbb{R}^{N\times 3}$. This publicly available dataset can be accessed at: https://doi.org/10.5281/zenodo.10998285.
@@ V Augmentations
We use a series of augmentations during the training of the contrastive learning models, as well as when training a baseline classifier with augmentations, which refers to a standard supervised classifier trained with data augmentations applied to the input samples during training. This ensures that the classifier is exposed to a range of transformations similar to those used in the contrastive learning setup, allowing a direct comparison of their effects on robustness to domain shifts. We apply the same augmentations to both datasets. These datasets were augmented using a series of transformations designed to introduce spatial and energy variations, simulating realistic conditions and improving model robustness. During training, for both the contrastive models and the classifier with augmentations, we randomly sample with replacement a series of three augmentations from a full set of augmentation transformations. The augmentations applied include: a rotation transformation that randomly rotates the spatial coordinates; a masking operation that randomly drops a fraction of the input points, effectively simulating partial data loss; an energy scaling augmentation that introduces voxel-wise energy shifts by scaling the feature values with random noise sampled from a normal distribution; a spatial translation transformation that displaces the input points by a random offset proportional to a predefined scale factor; and the identity transformation. These augmentations ensured compatibility with batched processing and scalable data handling.
We also experimented with augmentation via re-simulation of the detector response for the LArTPC dataset, similar to approaches explored in high-energy physics [30]. We re-simulated events at the larnd-sim level, altering parameters related to the electric field, transverse and longitudinal diffusion and electron lifetime. However, when using these augmentations with realistic simulation parameter ranges, we found the augmentations to be too weak. In line with previous work (making the case for using strong augmentations) [9], our trained contrastive model using purely re-simulation was unable to produce any useful features for downstream tasks. We also tried using re-simulation augmentations to create additional augmentations to complement the ones described in the previous paragraph. This resulted in marginal gains even when evaluated on samples that had undergone re-simulation with the same altered parameters as used in the augmentations. Given the significant computational requirement, we advocate against the use of re-simulation for augmentations in this use case.
@@ VI Experimental results
Figure 3: Bar chart of test accuracy of our methods - unsupervised and supervised contrastive learning, compared to a DANN and a classifier with training data augmentations. Nominal indicates the test accuracy of the models using the same detector simulation parameters used in training (i.e., no distribution shift). During testing, these methods are evaluated on various data throws - shifts in detector simulation parameters chosen randomly as shown in Table 1.
Figure 4: Top: Accuracy of our proposed methods - supervised and unsupervised contrastive learning compared to a classifier trained with augmentations on the scintillator detector dataset. Bottom: Relative change from nominal (training) level of cross-talk to the current (testing) cross-talk level. Nominal cross-talk is the level used to train the classifier and fine-tune the contrastive learning methods. The models are evaluated on cross-talk increments of 0.05-0.10, ranging from no cross-talk to the maximum level of 1.00.
@@
We evaluated the robustness of four different approaches: a classifier with augmentations, unsupervised contrastive learning, supervised contrastive learning, and DANN. Under nominal conditions, the classifier with augmentations achieved the highest performance at 87.3%, followed closely by our proposed method based on supervised contrastive learning (87.0%). The unsupervised contrastive learning approach resulted in a 3./drop in accuracy due to its inability to utilise the label information present in the dataset during the pre-training phase. The DANN approach led to a substantial performance decline, reaching only 74%, primarily due to the challenges in the adversarial training and the limited availability of diverse training examples for each throw.
@@
In Figure 4, we summarise the results of our proposed method on the scintillator detector dataset. Based on supervised and unsupervised contrastive learning, our methods are compared against a classifier using the same architecture. In Appendix A, we also compare the classifier without augmentations and a DANN.
The classifier is trained with the same augmentations applied in the contrastive learning pre-training setup. We generate datasets with different levels of cross-talk by masking out fractions of the detector response induced by cross-talk in the simulation. We train separate models for each cross-talk level and evaluate these models across all cross-talk levels. The supervised contrastive learning model consistently outperforms all other models except when the models are trained without any cross-talk. The unsupervised contrastive model exhibits almost no degradation as the data domain shifts (changes in cross-talk) during test time. Both contrastive setups maintain superior accuracy (compared to the classifier) as the cross-talk level varies, highlighting that the extracted features depend more on the fundamental physical properties of the particle interactions generating the events. In contrast, the classifier-based approach appears less robust, potentially indicating that its features are more sensitive to specific properties of the detector simulation.
These results strongly suggest that contrastive learning approaches provide enhanced stability against detector simulation variations compared to traditional deep learning classification methods. Looking at Equation 2, in this study, we have only considered setting $\alpha$, the factor controlling the weight of same-image versus same-class positives, to 0 or 1, corresponding to fully unsupervised and supervised contrastive learning. Depending on the task, other values may be considered, allowing us to prioritise robustness versus accuracy in the training data domain. The study of both unsupervised and supervised approaches indicates that the contrastive learning approach, rather than only its facilitation of pre-training on an unlabelled target distribution as in the unsupervised case, results in representations robust to distribution shifts. This increased stability is particularly valuable in high-energy physics and related domains, where minor variations in detector conditions can significantly impact model performance. By leveraging contrastive learning, we can ensure that learned representations generalise more effectively across varying conditions, thereby reducing the risk of model performance degradation due to distribution shifts.
@@ Acknowledgements.
Appendix A Results for all Models
Figure 5: Top: Accuracy of our proposed methods - supervised and unsupervised contrastive learning on the scintillator detector dataset. Our methods are compared to a DANN, and classifiers are trained with and without augmentations. Bottom: Relative change from nominal (training) level of cross-talk to the current (testing) cross-talk level. Nominal cross-talk is the level used to train the classifiers, fine-tune the contrastive learning methods, and provide the labelled source data for the DANN. The models are evaluated on cross-talk increments of 0.05-0.10, ranging from no cross-talk to the maximum level of 1.00.
In Figure 5, we present the results for all trained models on the scintillator detector dataset. Here, we include all models discussed in the main text in Figure 4 with the addition of a classifier trained without any augmentations and a Domain-Adversarial Neural Network (DANN). The DANN model demonstrates limited accuracy and exhibits substantial performance variability. In contrast, the contrastive learning-based models consistently achieve superior performance and demonstrate enhanced robustness to variations encountered during test time.